A tennis ball rolls off the edge of a table top 0.75 m above the floor and strikes the floor at a point 1.40 m horizontally from the edge of the table. Ignore air resistance. a). Find the time of flight. b). Find the magnitude of the initial velocity. c). Find the magnitude and direction of the velocity of the ball just before it strikes the floor


1
Expert's answer
2021-09-18T15:00:35-0400

a) "h=gt^2/2\to t=\sqrt{2h/g}=\sqrt{2\cdot0.75/9.81}=0.39\ (s)"


b) "v_i=l/t=1.4/0.39=3.6\ (m/s)"


c) "v_x=3.6\ (m/s)"

"v_y=gt=9.81\cdot0.39=3.8\ (m/s)"


"v=\sqrt{3.6^2+3.8^2}=5.2\ (m/s)"


"\tan\alpha=\frac{v_y}{v_x}=\frac{3.8}{3.6}=1.0555\to \alpha=46.5°" (below the horizontal)